OpenStudy (anonymous):

FAN+MEDAL************ y=x y=4x-6

TheSmartOne (thesmartone):

plug in equation 1 into equation 2

TheSmartOne (thesmartone):

so you get y = 4y - 6 solve for y

OpenStudy (anonymous):

-6/4y?

TheSmartOne (thesmartone):

??

TheSmartOne (thesmartone):

y = 4y - 6 add 6 on both sides subtract y on both sides

OpenStudy (anonymous):

should i divide the -6 by 4x?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

6y=4x+0

TheSmartOne (thesmartone):

noooo

TheSmartOne (thesmartone):

y = 4y - 6 add 6 on both sides y + 6 = 4y - 6 + 6 Simplify it more y + 6 = 4y Now you do this: subtract y on both sides

OpenStudy (anonymous):

6=4

OpenStudy (anonymous):



TheSmartOne (thesmartone):

4y - y = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

3y

TheSmartOne (thesmartone):

6 = 3y y = ?

OpenStudy (anonymous):

2

TheSmartOne (thesmartone):

ok, now we plug in y = 2 into y = x y = x 2 = x what is x = ? :P

OpenStudy (anonymous):

2

TheSmartOne (thesmartone):

Correct, so the solution to your system of equations is (2,2)

OpenStudy (anonymous):

Thank you z:)

TheSmartOne (thesmartone):

Anytime! :)